OpenStudy (anonymous):

Use the distributive property to simplify the expression below: The quantity of negative 4 times x minus 16 all divided by negative 4

OpenStudy (unklerhaukus):

\[\frac{-4x-16}{-4}=(-4x-16)\div-4\] The distributive property of an operator \(\star\), says that: \[(a+b)\star c=(a\star c)+(b\star c)\] so \[(-4x-16)\div-4=(-4x\div-4)+(-16\div-4)\\\qquad\qquad\qquad\qquad=\frac{-4x}{-4}+\frac{-16}{-4}\]

OpenStudy (unklerhaukus):

can you simplify further from here @cosmicchild?